HETAV PATEL

Grade 11,

how many lattice points are their in one unit cell of each of the following lattice? 1)face centered cubic 2)face centered tetragonal 3)body centered

how many lattice points are their in one unit cell of each of the following lattice?
1)face centered cubic
2)face centered tetragonal
3)body centered

Grade:12

5 Answers

BOLLU SRINIVAS
8 years ago
(i) Number of corner atoms per unit cell= 8 corners atoms X1/8 atom per unit cell= 8 X 1/8 = 1 atomNumber of face centred atoms per unit cell= 6 face centred atoms X 1/2 atom per unit cell= 6 X1/2 = 3 atoms
BOLLU SRINIVAS
8 years ago
 Total number of atoms or lattice points = 1 + 3 = 4.
(ii) No. of lattice points in Face-centred tetragonal = 4.
(iii) In bcc unit cell, number of corner atoms per unit cell
= 8 corners X 1/8 per corner atom.
= 8 X 1/8 = 1 atom
Number of atoms at body centre = 1 X 1 = 1 atom
Total number of atoms (lattice points) = 1 + 1 = 2.
 
Priyanshu
5 years ago
Lattice point are there in 1 unit cell in each of the following letters are as follows:-(1). Face centered cubic have 4(2). Face centred tetragonal have 4(3). Body centred have 2
Gitanjali Rout
5 years ago
i) Number of corner atoms per unit cell= 8 corners atoms X1/8 atom per unit cell= 8 X 1/8 = 1 atomNumber of face centred atoms per unit cell= 6 face centred atoms X 1/2 atom per unit cell= 6 X1/2 = 3 atoms
Gitanjali Rout
5 years ago
otal number of atoms or lattice points = 1 + 3 = 4.
(ii) No. of lattice points in Face-centred tetragonal = 4.
(iii) In bcc unit cell, number of corner atoms per unit cell
= 8 corners X 1/8 per corner atom.
= 8 X 1/8 = 1 atom
Number of atoms at body centre = 1 X 1 = 1 atom